Calculate the specific heat of a 34 g sample of metal if it absorbs 7500 Jules of he and his temperature changes 24°C


1
Expert's answer
2021-01-15T06:48:33-0500

c = Q/m/"\Delta"T = (7500 J)/(34 g)/(24 °C) = 9.19118 J/g/°C = 9191.18 J/kg/°C